A long solenoid (1000 turns/m) carries a current of 25 mA and has an inside radius of 2.0 cm. A long wire which is parallel to and 4.0 cm from the axis of the solenoid carries a current of 6.0 A. Determine the magnitude of the magnetic field at a point on the axis of the solenoid.
1
Expert's answer
2020-03-31T09:37:00-0400
"B=\frac{\mu_0 I}{2\pi r}"

"B=\frac{4\pi\cdot 10^{-7} 6}{2\pi (0.04)}=3.0\cdot 10^{-5}\ T"